# Failure of uniform laws of large numbers for subdifferentials and beyond

### v1

@@ 2.2 Random Convex Functions
Section 2.1 shows that the uniform laws in (1) and (2) with $r=0$ cannot hold for random Lipschitz functions. One might speculate that Lipschitz functions are somewhat “wild” and the situation would be better for other nonsmooth functions. A natural class to examine next is that of random convex functions, since convexity is widely regarded as favorable. The following uniform law for $\varepsilon$-subdifferentials is classical in the literature.
Theorem 2 (cf. [18, Proposition 3.4] and [20, Theorem 7.56]).
Let nonempty $X\subset\mathbb{R}^{d}$ be compact, $\Xi\subset\mathbb{R}^{m}$, and $\bm{\xi}^{1},\bm{\xi}^{2},\ldots$ be $\Xi$-valued iid random variables on the complete probability space $(\Omega,\mathcal{F},\mathbb{P})$. Let $f:\Xi\times\mathbb{R}^{d}\to\mathbb{R}$ be a Carathéodory function such that $f(\xi,\cdot)$ is convex for any $\xi\in\Xi$. Moreover, assume that $|\mathbb{E}[f(\bm{\xi},x)]|<\infty$ for any $x\in\mathbb{R}^{d}$. Then, for any fixed $\varepsilon>0$, $\mathbb{P}$-almost surely, one has
@@
Theorem 2 differs in several ways from the enlarged uniform laws in (1) and (2). First, there is no Lipschitz assumption in Theorem 2. Second, for a convex $h$, unlike the $r$-enlargement $\cup_{y\in\mathbb{B}(x,r)}\partial h(y)$, the $\varepsilon$-subdifferential $\partial^{\varepsilon}h(x)$ is intrinsically a global notion by definition. Third, the mapping $x\mapsto\partial^{\varepsilon}h(x)$ is continuous with respect to the Hausdorff metric, whereas the mapping $x\mapsto\cup_{y\in\mathbb{B}(x,r)}\partial h(\xi,y)$ is not. Moreover, as emphasized in [20, p. 382], for random convex function $f:\Xi\times\mathbb{R}^{d}\to\mathbb{R}$ and $\xi^{1},\ldots,\xi^{\nu}\in\Xi$, the subdifferential sum rule
$\partial^{\varepsilon}\big(\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}f(\xi^{i},\cdot)\big)(x)=\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}\partial_{x}^{\varepsilon}f(\xi^{i},x)$
holds for $\varepsilon=0$, but fails for $\varepsilon>0$ and $\nu>1$.
Therefore, for uniform laws of the form (1) and/or (2) (and, when both are considered, equivalently (6)) for random convex functions, it is natural to ask whether one can take $r=0$ (or $\varepsilon=0$ in (6)). We next see that the answer to this question crucially depends on the dimension $d$ of $x$.
@@
For a measurable space $(\Xi,\mathcal{A})$, let $\bm{\xi}^{1},\bm{\xi}^{2},\ldots$ be $\Xi$-valued iid random variables on the complete probability space $(\Omega,\mathcal{F},\mathbb{P})$. Let $f:\Xi\times\mathbb{R}\to\mathbb{R}$ be a Carathéodory function such that $f(\xi,\cdot)$ is convex and $L(\xi)$-Lipschitz for any $\xi\in\Xi$ with measurable $L$ and $\mathbb{E}[L(\bm{\xi})]<\infty$. Assume that $|\mathbb{E}[f(\bm{\xi},x)]|<\infty$ for any $x\in\mathbb{R}$. Then, $\mathbb{P}$-almost surely, one has
$\lim_{\nu\to\infty}\sup_{x\in\mathbb{R}}{d\kern-1.49994ptl}\Big(\partial\mathbb{E}[f(\bm{\xi},\cdot)](x),\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}\partial_{x}f(\bm{\xi}^{i},x)\Big)=0.$
In sharp contrast to the univariate random Lipschitz case, convexity allows us to take $r=0$ in both (1) and (2), even without any compactness assumption. The uniform law in Proposition 1 holds on the entire real line. In spirit, it is similar to the classical Glivenko–Cantelli theorem and, even when restricted to a compact $X$, already covers cases that were previously unknown. For illustration, let $X=[0,3]$, and let $\bm{\xi}^{\nu}=(\bm{\xi}^{\nu}_{1},\bm{\xi}^{\nu}_{2})$ be iid random variables with $\bm{\xi}^{\nu}_{1}$ uniformly distributed on $[0,1]$ and $\bm{\xi}^{\nu}_{2}=2$ almost surely. Define $f(\xi,x)=\max\{x-\xi_{1},0\}+\max\{\xi_{2}-x,0\}$. It is easy to see that $x\mapsto\mathbb{E}[f(\bm{\xi},x)]$ is not differentiable (it has a kink at $x=2$), and a computation similar to [12, Example 3.4] shows that the range of $\partial_{x}f$ is not essentially separable. Moreover, one can show that $f$ cannot be written in the univariate convex-composite form considered in [15, Section 4]. Hence, a subdifferential uniform law for $f$ cannot be deduced from the results in [17], [21], or [15]. However, Proposition 1 applies.
For the univariate case ($d=1$), we thus have obtained reasonably satisfactory positive results. What about higher dimensions?
@@
Let $X=\{0\}\times[0,1]\subset\mathbb{R}^{2}$, and $\bm{\xi}^{1},\bm{\xi}^{2},\ldots$ be iid random variables on the complete probability space $(\Omega,\mathcal{F},\mathbb{P})$, each uniformly distributed on $\Xi=[0,1]$. There exist a Carathéodory $g:\Xi\times\mathbb{R}^{2}\to\mathbb{R}$ and $\delta^{\nu}\in(0,1]\downarrow 0$ such that for
@@
In Theorem 3, we show that even in dimension $d=2$ there exists a random convex function $f:\Xi\times\mathbb{R}^{2}\to\mathbb{R}$ for which a negative result analogous to Theorem 1 holds. Therefore, in the same vein as Corollary 1, one cannot take $r=0$ in (1) and (2); equivalently, a uniform law with $\varepsilon=0$ in (6) cannot hold. The same result holds for any $d\geq 2$ simply by padding with zero coordinates.
However, Theorem 3 says more than this, since the random convex $f$ there is highly structured. Let us make the following observations:
  • In contrast to the construction used to prove Theorem 1 (see Section 3.2), which involves countably infinitely many pieces [^1] Here, by the number of “pieces,” we mean the smallest number of connected regions partitioning the domain such that the function is smooth on the interior of each region. (see Figure 1), the function $f$ in Theorem 3 is a random convex function with only two pieces (see Figure 2). Indeed, the only source of nonsmoothness in the construction is the function $t\mapsto\max\{t,0\}$. Moreover, by [20, Theorem 7.52], we have
@@
    for the function $g$ emerging from Theorem 3, $\mathbb{P}$-almost surely, with $X$ being any nonempty compact set. Thus, each smooth component appearing in the definition of $f$ satisfies a uniform law on its own. However, when they are combined through this (arguably) simplest nonsmooth operation $t\mapsto\max\{t,0\}$, the uniform law fails. This phenomenon is driven by the oscillating boundary between the connected sets $\{x\mid g(\xi,x)>0\}$ and $\{x\mid g(\xi,x)<0\}$, whose intersection with $X=\{0\}\times[0,1]$ gives rise to countably infinitely many kinks in the univariate function $t\mapsto\max\{g(\xi,(0,t)),0\}$.
  • The negative results in Theorem 3 hold for any function class that contains $f$. An interesting example is a subclass of the random convex-composite functions studied in [15, Section 4]. For $\kappa<\infty$, consider a convex, piecewise affine, $\kappa$-Lipschitz $h:\mathbb{R}\to\mathbb{R}$ and a Carathéodory $c:\Xi\times\mathbb{R}^{d}\to\mathbb{R}$ such that $c(\xi,\cdot)$ is $\kappa$-Lipschitz and $\kappa$-smooth for each $\xi$. Since $\kappa$ is finite and $h$ is piecewise affine, the constants in [15, Assumptions C.1–C.3] are distribution-free. By [15, Theorem 5], for large $\nu$, with (outer) probability at least $1-\delta$, one has
    $\sup_{x\in B}{d\kern-1.49994ptl}\Big(\partial\mathbb{E}[(h\circ c)(\bm{\xi},\cdot)](x),\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}\partial_{x}(h\circ c)(\bm{\xi}^{i},x)\Big)\leq O\bigg(\sqrt{\frac{d+\textsf{VC}(\mathcal{C})\log\nu+\log(1/\delta)}{\nu}}\bigg),$
    where $B$ is an open Euclidean ball, $\mathcal{C}=\{\{\xi\mid c(\xi,x)\geq t\}\mid x\in\mathbb{R}^{d},t\in\mathbb{R}\}$, and $\textsf{VC}(\mathcal{C})$ denotes the VC dimension of $\mathcal{C}$; see [22, Section 2.6.1]. This upper bound is also distribution-free, but is non-trivial only if $\textsf{VC}(\mathcal{C})<\infty$; in other words, the function class $\{c(\cdot,x)\mid x\in\mathbb{R}^{d}\}$ is VC-major; see [22, Section 2.6.4, Exercise 2.6.13]. The VC-type assumptions are usually employed to obtain distribution-free results [22, Section 2.8.1] and may not be necessary for universal Glivenko–Cantelli-type results, let alone $\mathbb{P}\circ(\bm{\xi}^{1})^{-1}$-Glivenko–Cantelli-type results such as [20, Theorem 7.52]. Hence, it is natural to ask whether the VC-major assumption here can be dispensed with when uniformity over the underlying distribution and an explicit convergence rate are not required. Consider the convex piecewise affine $h=\max\{\cdot,0\}$ and let $c=g$, where $g$ is as in Theorem 3. We can write the function $f$ in Theorem 3 as
@@
  • Another example is median regression. Let $\bm{\xi}^{\nu}=(\bm{\xi}_{1}^{\nu},\bm{\xi}_{2}^{\nu})$ be iid $\Xi\times[-1,1]$-valued random variables with $\Xi\subset\mathbb{R}^{m}$, and let $\psi:\Xi\times\mathbb{R}^{d}\to\mathbb{R}$ be a Carathéodory function such that $\psi(\xi_{1},\cdot)$ is Lipschitz and smooth for each $\xi_{1}\in\Xi$. Median regression corresponds to the random function $f(\xi,x)=|\xi_{2}-\psi(\xi_{1},x)|$. It is easy to see that the construction in Theorem 3 can be rewritten in this form, since $f(\xi,x)=\max\{2\xi_{2}-2\psi(\xi_{1},x),0\}-\xi_{2}+\psi(\xi_{1},x)$ with the last term being handled by [20, Theorem 7.52]. Therefore, in general, uniform laws fail for the subdifferential of $f$ when $d>1$.
@@
Under the assumptions of Theorem 3, let $X$, $\Xi$, $\{\bm{\xi}^{\nu}\}_{\nu}$, $f$, and $\{\delta^{\nu}\}_{\nu}$ be constructed there. For any sequence $\{\varepsilon^{\nu}\in[0,(\delta^{\nu})^{2}]\}_{\nu}$, $\mathbb{P}$-almost surely, one has
@@
For any $x^{\nu}\in D^{\nu}$ in Theorem 3, let $g_{1}^{\nu}\in\partial^{\varepsilon^{\nu}}\mathbb{E}[f(\bm{\xi},\cdot)](x^{\nu})$ and $g_{2}^{\nu}\in\partial^{\varepsilon^{\nu}}\big(\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}f(\bm{\xi}^{i},\cdot)\big)(x^{\nu})$. By Brøndsted–Rockafellar [2, p. 608] and $\varepsilon^{\nu}\in[0,(\delta^{\nu})^{2}]$, we get $\|g_{1}^{\nu}-\mathbb{E}[\nabla_{x}f(\bm{\xi},y^{\nu}_{1})]\|\leq\delta^{\nu}$ and $\|g_{2}^{\nu}-\tfrac{1}{\nu}\sum_{i=1}^{\nu}\nabla_{x}f(\bm{\xi}^{i},y^{\nu}_{2})\|\leq\delta^{\nu}$, where $y_{1}^{\nu},y_{2}^{\nu}\in\mathbb{B}(x^{\nu},\delta^{\nu})$. Then, $\|g_{1}^{\nu}-g_{2}^{\nu}\|\geq\|\mathbb{E}[\nabla_{x}f(\bm{\xi},y^{\nu}_{1})]-\tfrac{1}{\nu}\sum_{i=1}^{\nu}\nabla_{x}f(\bm{\xi}^{i},y^{\nu}_{2})\|-2\delta^{\nu}$. The proof completes by invoking Theorem 3 and $\delta^{\nu}\downarrow 0$. ∎

### v2

@@ 2.2 Random Convex Functions
Section 2.1 shows that the uniform laws in (1) and (2) with $r=0$ cannot hold for random Lipschitz functions. One might speculate that Lipschitz functions are somewhat “wild” and the situation would be better for other nonsmooth functions. A natural class to examine next is that of random convex functions, since convexity is widely regarded as favorable. The following uniform law for $\varepsilon$-subdifferentials is classical.
Theorem 2 (cf. [23, Proposition 3.4] and [25, Theorem 7.56]).
Let nonempty $X\subset\mathbb{R}^{d}$ be compact, $\Xi\subset\mathbb{R}^{m}$, and $\bm{\xi},\bm{\xi}^{1},\bm{\xi}^{2},\ldots$ be $\Xi$-valued iid random variables on the complete probability space $(\Omega,\mathcal{F},\mathbb{P})$. Let $f:\Xi\times\mathbb{R}^{d}\to\mathbb{R}$ be a Carathéodory function such that $f(\xi,\cdot)$ is convex for any $\xi\in\Xi$. Assume that $|\mathbb{E}[f(\bm{\xi},x)]|<\infty$ for any $x\in\mathbb{R}^{d}$. Then, for any fixed $\varepsilon>0$, $\mathbb{P}$-a.s., one has
@@
Theorem 2 differs in several ways from the enlarged uniform laws in (1) and (2). First, there is no Lipschitz assumption in Theorem 2. Second, for a convex real-valued $h$ and a point $x$, unlike the $r$-enlargement $\cup_{y\in\mathbb{B}(x,r)}\partial h(y)$, the $\varepsilon$-subdifferential $\partial^{\varepsilon}h(x)$ is intrinsically a global notion by definition. Third, the mapping $x\mapsto\partial^{\varepsilon}h(x)$ with $\varepsilon>0$ is continuous with respect to the Hausdorff metric, whereas the mapping $x\mapsto\cup_{y\in\mathbb{B}(x,r)}\partial h(y)$ may fail to be. Moreover, as emphasized in [25, p. 382], for $\xi^{1},\ldots,\xi^{\nu}\in\Xi$ and a function $f:\Xi\times\mathbb{R}^{d}\to\mathbb{R}$ such that $f(\xi,\cdot)$ is convex for every $\xi$, the $\varepsilon$-subdifferential sum rule
$\partial^{\varepsilon}\big(\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}f(\xi^{i},\cdot)\big)(x)=\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}\partial_{x}^{\varepsilon}f(\xi^{i},x),$
where $\partial_{x}^{\varepsilon}f(\xi^{i},x)=\partial^{\varepsilon}(f(\xi^{i},\cdot))(x)$, holds for $\varepsilon=0$ but can fail for $\varepsilon>0$ and $\nu>1$.
Therefore, for uniform laws of the form (1) and/or (2) (also (6)) for random convex functions, it is natural to ask whether one can take $r=0$ (equivalently, $\varepsilon=0$ in (6)). We next see that the answer to this question crucially depends on the dimension $d$ of $x$.
@@
For a Borel measurable $\Xi\subset\mathbb{R}^{m}$, let $\bm{\xi},\bm{\xi}^{1},\bm{\xi}^{2},\ldots$ be $\Xi$-valued iid random variables on the complete probability space $(\Omega,\mathcal{F},\mathbb{P})$. Let $f:\Xi\times\mathbb{R}\to\mathbb{R}$ be a Carathéodory function such that $f(\xi,\cdot)$ is convex and $L(\xi)$-Lipschitz for any $\xi\in\Xi$ with Borel measurable $L$ and $\mathbb{E}[L(\bm{\xi})]<\infty$. Assume that $|\mathbb{E}[f(\bm{\xi},x)]|<\infty$ for any $x\in\mathbb{R}$. Then, $\mathbb{P}$-a.s., one has
$\lim_{\nu\to\infty}\sup_{x\in\mathbb{R}}{d\kern-1.49994ptl}\Big(\mathbb{E}[\partial_{x}f(\bm{\xi},x)],\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}\partial_{x}f(\bm{\xi}^{i},x)\Big)=0.$
In sharp contrast to the univariate random Lipschitz case, convexity allows us to take $r=0$ in both (1) and (2), even without any compactness assumption. The uniform law in Proposition 1 holds on the entire real line. In spirit, it is similar to the classical Glivenko–Cantelli theorem and, even when restricted to a compact $X$, already covers cases that were previously unknown. For illustration, let $X=[0,3]$, and let $\bm{\xi}=(\bm{\xi}_{1},\bm{\xi}_{2})$, where $\bm{\xi}_{1}$ and $\bm{\xi}_{2}$ are independent random variables, with $\bm{\xi}_{1}$ uniformly distributed on $[0,1]$ and $\bm{\xi}_{2}=2$ almost surely. Define $f(\xi,x)=\max\{x-\xi_{1},0\}+\max\{\xi_{2}-x,0\}$. It is easy to see that $x\mapsto\mathbb{E}[f(\bm{\xi},x)]$ is not differentiable (it has a kink at $x=2$), and a computation similar to [16, Example 3.4] shows that the range of $\partial_{x}f$ is not separable. Moreover, one can show that $f$ cannot be written in the univariate convex-composite form considered in [20, Section 4]. Hence, a subdifferential uniform law for $f$ cannot be deduced from the results in [22, 24, 26, 30, 20]. However, Proposition 1 applies.
@@
Let $X=\{0\}\times[0,1]\subset\mathbb{R}^{2}$, and $\bm{\xi},\bm{\xi}^{1},\bm{\xi}^{2},\ldots$ be iid random variables on the complete probability space $(\Omega,\mathcal{F},\mathbb{P})$, each uniformly distributed on $\Xi=[0,1]$. There exist a Carathéodory function $g:\Xi\times\mathbb{R}^{2}\to\mathbb{R}$ and $\delta^{\nu}\in(0,1]\downarrow 0$ such that for
@@
In Theorem 3, we show that even in dimension $d=2$ there exists a random convex function $f:\Xi\times\mathbb{R}^{2}\to\mathbb{R}$ for which a negative result analogous to Theorem 1 holds. Therefore, in the same vein as Corollary 1, one cannot take $r=0$ in (1) and (2); equivalently, a uniform law with $\varepsilon=0$ in (6) cannot hold. The same lower bound extends to all $d\geq 2$ by padding with zero coordinates, and to any compact superset of $X$. Thus, restricting to a lower-dimensional set $X$ strengthens, rather than weakens, the negative result.
However, Theorem 3 says more than these, since the random convex $f$ there is highly structured. Let us make the following observations:
  • In contrast to the construction used to prove Theorem 1 (see Section 3.2), which involves countably infinitely many pieces [^1] Here, by the number of “pieces,” we mean the smallest number of connected regions partitioning the domain such that the function is smooth on the interior of each region. (see Figure 1), the function $f$ in Theorem 3 is a random convex function with only two pieces (see Figure 2). Indeed, the only source of nonsmoothness in the construction is the function $t\mapsto\max\{t,0\}$. Moreover, by [25, Theorem 7.52], we have
@@
    for the function $g$ emerging from Theorem 3, $\mathbb{P}$-a.s., with $X$ being any nonempty compact set. Thus, each smooth component appearing in the definition of $f$ satisfies a uniform law on its own. However, when they are combined through this (arguably) simplest nonsmooth operation $t\mapsto\max\{t,0\}$, the uniform law fails. This phenomenon is driven by the oscillating boundary between the connected sets $\{x\mid g(\xi,x)>0\}$ and $\{x\mid g(\xi,x)<0\}$, whose intersection with $X=\{0\}\times[0,1]$ gives rise to countably infinitely many kinks in the univariate function $t\mapsto\max\{g(\xi,(0,t)),0\}$, leading to an oscillating “active set.”
  • The negative results in Theorem 3 hold for any function class that contains $f$. An interesting example is a subclass of the random convex-composite functions studied in [20, Section 4]. For $\kappa<\infty$, consider a convex, piecewise affine, $\kappa$-Lipschitz $h:\mathbb{R}\to\mathbb{R}$ and a Carathéodory $c:\Xi\times\mathbb{R}^{d}\to\mathbb{R}$ such that $c(\xi,\cdot)$ is $\kappa$-Lipschitz and $\kappa$-smooth for each $\xi$. Since $\kappa$ is finite and $h$ is piecewise affine, the constants in [20, Assumptions C.1–C.3] are distribution-free. By [20, Theorem 5], for large $\nu$, with (outer) probability at least $1-\delta$, one has
    $\sup_{x\in B}{d\kern-1.49994ptl}\Big(\mathbb{E}[\partial_{x}(h\circ c)(\bm{\xi},x)],\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}\partial_{x}(h\circ c)(\bm{\xi}^{i},x)\Big)\leq O\bigg(\sqrt{\frac{d+\textsf{VC}(\mathcal{C})\log\nu+\log(1/\delta)}{\nu}}\bigg),$
    where $B$ is an open Euclidean ball, $\mathcal{C}=\{\{\xi\mid c(\xi,x)\geq t\}\mid x\in\mathbb{R}^{d},t\in\mathbb{R}\}$, and $\textsf{VC}(\mathcal{C})$ denotes the VC dimension of $\mathcal{C}$; see [28, Section 2.6.1]. This upper bound is also distribution-free, but is non-trivial only if $\textsf{VC}(\mathcal{C})<\infty$; in other words, the function class $\{c(\cdot,x)\mid x\in\mathbb{R}^{d}\}$ is VC-major; see [28, Section 2.6.4, Exercise 2.6.13]. The VC-type assumptions are usually employed to obtain distribution-free results [28, Section 2.8.1] and may not be necessary for universal Glivenko–Cantelli-type results, let alone $\mathbb{P}\circ\bm{\xi}^{-1}$-Glivenko–Cantelli-type results such as [25, Theorem 7.52]. Hence, it is natural to ask whether the VC-major assumption here can be dispensed with when uniformity over the underlying distributions and an explicit convergence rate are not required. Consider the convex piecewise affine $h=\max\{\cdot,0\}$ and let $c=g$, where $g$ is as in Theorem 3. We can write the function $f$ in Theorem 3 as
@@
  • Another example is median regression. Let $\bm{\xi}=(\bm{\xi}_{1},\bm{\xi}_{2})$, where $\bm{\xi}_{1}$ and $\bm{\xi}_{2}$ are $\Xi$-valued and [1,1][-1,1]-valued random variables, respectively, with $\Xi\subset\mathbb{R}^{m}$. Let $\psi:\Xi\times\mathbb{R}^{d}\to\mathbb{R}$ be a Carathéodory function such that $\psi(\xi_{1},\cdot)$ is Lipschitz and smooth for each $\xi_{1}\in\Xi$. Median regression corresponds to the function $f(\xi,x)=|\xi_{2}-\psi(\xi_{1},x)|$. It is easy to see that the construction in Theorem 3 can be rewritten in this form, since $f(\xi,x)=\max\{2\xi_{2}-2\psi(\xi_{1},x),0\}-\xi_{2}+\psi(\xi_{1},x)$ with the last term being handled by [25, Theorem 7.52]. Therefore, in general, uniform laws fail for the subdifferential of $\mathbb{E}[f(\bm{\xi},\cdot)]$ when $d>1$.
@@
Under the assumptions of Theorem 3, let $X$, $\Xi$, $\{\bm{\xi},\bm{\xi}^{\nu}\}_{\nu}$, $f$, and $\{\delta^{\nu}\}_{\nu}$ be constructed there. For any sequence $\{\varepsilon^{\nu}\in[0,(\delta^{\nu})^{2}]\}_{\nu}$, $\mathbb{P}$-a.s., one has
@@
For any $x^{\nu}\in D^{\nu}$ in Theorem 3, let $g_{1}^{\nu}\in\partial^{\varepsilon^{\nu}}\mathbb{E}[f(\bm{\xi},\cdot)](x^{\nu})$ and $g_{2}^{\nu}\in\partial^{\varepsilon^{\nu}}\big(\tfrac{1}{\nu}\mathop{\sum}\nolimits_{i=1}^{\nu}f(\bm{\xi}^{i},\cdot)\big)(x^{\nu})$ be arbitrary. By Brøndsted–Rockafellar [5, p. 608] and $\varepsilon^{\nu}\in[0,(\delta^{\nu})^{2}]$, we get $\|g_{1}^{\nu}-\mathbb{E}[\nabla_{x}f(\bm{\xi},y^{\nu}_{1})]\|\leq\delta^{\nu}$ and $\|g_{2}^{\nu}-\tfrac{1}{\nu}\sum_{i=1}^{\nu}\nabla_{x}f(\bm{\xi}^{i},y^{\nu}_{2})\|\leq\delta^{\nu}$, where $y_{1}^{\nu},y_{2}^{\nu}\in\mathbb{B}(x^{\nu},\delta^{\nu})$. Then, $\|g_{1}^{\nu}-g_{2}^{\nu}\|\geq\|\mathbb{E}[\nabla_{x}f(\bm{\xi},y^{\nu}_{1})]-\tfrac{1}{\nu}\sum_{i=1}^{\nu}\nabla_{x}f(\bm{\xi}^{i},y^{\nu}_{2})\|-2\delta^{\nu}$. This completes the proof by invoking Theorem 3 and $\delta^{\nu}\downarrow 0$. ∎